OpenStudy (anonymous):

Convert 48 miles per hour to kilometers per minute. Show the ratio or series of ratios used to determine the answer.

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

conversion huh. 1 mile = 1.6 km 1 hr = 60 min so you have \[\large \frac{48 \;\; \text{miles}}{\text{hour}} \;\; \times \frac{1.6 \; \text{kilometers}}{1 \;\text{mile}} \times \frac{1 \; \text{hour}}{60 \; \text{minutes}}\] solve to get the conversion